OpenStudy (anonymous):

Show that the given value of c is a zero of P(x), and find all other zeros of P(x). P(x) = x3 − 4x2 − 3x + 12 , c = 4

OpenStudy (anonymous):

Evaluate P(c) to show that it is a zero then divide P(x) by (x-c) to get a quadratic equation from which you can get the other zeros.

OpenStudy (anonymous):

once I divided I got quotient x^2-3 and remainder 0. what do i do next?

OpenStudy (anonymous):

So now you know that you can write P(x)=(x-4)(x^2-3). So if we want P(x)=0 we want at least one of the factors to be zero. So your other zeros will arise from x^2-3=0. You can test these zeros in the original P(x) experession to be sure

OpenStudy (anonymous):

thanks (: